OpenStudy (anonymous):

At what point on the graph of y=(1/2)x^2-(3/2) is the tangent line parallel to the line 4x-8y=5?

OpenStudy (anonymous):

where the derivative is equal to \(\frac{1}{2}\)

OpenStudy (anonymous):

can u explain it plz? as in how u go to that point (finding 1/2)

zepdrix (zepdrix):

If you put 4x-8y=5 into slope-intercept form ( y=mx+b ), you can find the slope that you need to be PARALLEL, or in other words, EQUAL to. Remember, the derivative represents the SLOPE of your function at a particular point. We want to know when that SLOPE is the same as the slope of the linear function they gave us!

OpenStudy (anonymous):

the slope of line you are given is \(\frac{1}{2}\)

OpenStudy (anonymous):

if your curve is \(y=\frac{1}{2}x^2+\frac{3}{2}\) then the derivative is \(y'=x\)

OpenStudy (anonymous):

then since the derivative is just \(x\) you set \(x=\frac{1}{2}\) and there is nothing to solve